OpenStudy (anonymous):

sSolve the equation by a method of your choice: x2 - 6x + 5 = 0 If there are two solutions, separate the solutions by a comma.

OpenStudy (anonymous):

x^2-5x-x+5 =0 x(x-5)-1(x-5) = 0 (x-5)(x-1) = 0 either x-5 =0 or x-1 =0 that is x = 5,1

OpenStudy (anonymous):

so x=-5 x=-1

OpenStudy (mathstudent55):

x2 - 6x + 5 = 0 (x - 5)(x - 1) = 0 x - 5 = 0 or x - 1 = 0 x = 5 or x = 1 Final answer: 5, 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks